# Weak ergodicity breaking from supersymmetry in a fermionic kinetically constrained model

### v1

@@ Abstract
Supersymmetry provides a natural playground for the construction of kinetically constrained lattice fermion models showing weak ergodicity breaking. The supersymmetric algebra naturally induces non-ergodic dynamics, as we illustrate here by introducing a fermionic equivalent of the PXP model with an adjustable chemical potential. This model is closely related to the $\mathcal{N}=2$ supersymmetric $M_{1}$ model. Supersymmetry directly implies that the dynamics exhibit periodic revivals for specific initial states, including the $\mathbb{Z}_{2}$-ordered (every second site occupied) product state. These dynamics are reminiscent to those of the PXP model, a paradigmatic toy model in the field of quantum many-body scars. We draw a further parallel by uncovering quantum many-body scar-like eigenstates obeying sub-thermal entanglement scaling at energies given by (plus or minus) square roots of integers and relate these to special eigenstates of the $M_{1}$ model.
@@ I Introduction
Starting from the experimental observation of periodic revivals in the dynamics of programmable atom quantum simulators [1, 2], it is by now well-established that quantum many-body systems with (emergent) constraints can avoid thermalization up to time scales much longer than naively expected [3, 4, 5, 6]. The breakdown of ergodicity in such systems can often be related to a small number of highly non-thermal eigenstates known as quantum many-body scars [7, 8], in loose analogy to the single-particle quantum scars first observed by Heller in 1984 [9, 10, 11]. These eigenstates occur at finite energy density, thereby strongly violating the eigenstate thermalization hypothesis [12, 13, 14].
@@
Several mechanisms leading to weak ergodicity breaking (i.e., the partial breakdown of ergodicity) through the emergence of quantum many-body scars have been identified by now. Among these are the presence of an approximate hidden SU(2) algebra [15, 16, 17, 18], a spectrum generating algebra [19, 20, 21, 22, 23, 24], the proximity of the model to integrability [25, 26], or Bethe-ansatz integrability [27]. In this work, we focus on a less-studied mechanism that could induce weak ergodicity breaking in constrained quantum many-body systems: supersymmetry [28, 29, 30]. As we point out, supersymmetry imposes a non-ergodic structure on the eigenstates that can not be captured by ordinary conservation laws such as particle number or momentum conservation.
@@
The $M_{1}$ model provides a relatively well-studied example of a supersymmetric constrained lattice quantum many-body system [31, 32]. It describes short-range interacting spinless fermions on a chain (although it can naturally be defined on any graph), subject to the constraint that two neighboring sites can not be occupied simultaneously. The model is integrable by the Bethe ansatz [33, 34, 31, 32], and can be realized experimentally [35, 36]. Ref. [37] recently studied weak ergodicity breaking for the $M_{1}$ model on hypercubic lattice geometries. It was found that a number of highly non-thermal eigenstates, interpreted as quantum many-body scars, can be found thanks to the supersymmetry. In this work, we use the $M_{1}$ model on a chain as a starting point to construct a fermionic equivalent of the (hard-core bosonic) PXP model. The PXP model [38, 39, 7, 8] provides a paradigmatic toy model in studies on quantum many-body scars and weak ergodicity breaking [3, 4, 5, 6].
@@
The aim of this work is twofold: First, we show how the algebraic structure of supersymmetry can be used to construct kinetically constrained lattice Hamiltonians displaying weak ergodicity breaking (Sec. II). While the procedure holds more generally, we focus on the $M_{1}$ model for concreteness, for which the resulting Hamiltonian represents a fermionic version of the PXP model with an adjustable chemical potential:
@@
expressed in terms of fermionic creation/annihilation operators $c_{i}^{\dagger}$ and $c_{i}$, projectors $P_{i}=1-c_{i}^{\dagger}c_{i}$, and the total fermion number operator $F$. This model is introduced in detail in Sec. III. We show how the supersymmetry of the $M_{1}$ model leads to exact and approximate periodic revivals in the quench dynamics for specific initial states, including the $\mathbb{Z}_{2}$-ordered (every second site occupied) product state (Sec. IV). Periodic revivals for a small set of initial states, typically including the $\mathbb{Z}_{2}$-ordered product state, are often seen as a hallmark of quantum many-body scarring and have also been observed in the dynamics of the PXP model [7, 8].
Second, and of independent interest, we identify special eigenstates of the $M_{1}$ Hamiltonian on a chain, supporting a matrix product state (MPS) description with small bond dimension and a resulting sub-thermal entanglement scaling (Sec. V). These special eigenstates, occurring at energies given by plus or minus square roots of integers when considering the PXP-like fermion model, can be thought of as quantum many-body scars. Similar special eigenstates at energies given by plus or minus square roots of integers have similarly been observed for the PXP model [40, 41].
@@ II Supersymmetry
Supersymmetric quantum mechanics provides a versatile tool for the construction of constrained lattice fermion models showing weak ergodicity breaking. An $\mathcal{N}=2$ supersymmetric quantum model is described by a Hamiltonian of the form
@@
Observing that $[F,H_{S}]=0$ shows that the fermion number is a conserved quantity. In the following, we focus on Hamiltonians of the form
@@
where $\mu$ is a free parameter having the interpretation of a chemical potential. Such Hamiltonians have recently appeared in Ref. [42] in the context of central spin models. As one direct illustration of constraints that supersymmetry imposes on the dynamics of this model, we can consider the evolution of the total fermion number. While the total number of fermions $F$ is no longer conserved, the dynamics of the fermion number under the dynamics of $H$ are generically non-ergodic. Consider, for example, a state $|\psi\rangle$ with a fixed fermion number, i.e. $F|\psi\rangle=f|\psi\rangle$. This state is a superposition of eigenstates of the supersymmetric Hamiltonian (2) with $f$ fermions. The Hamiltonian (4) only couples these eigenstates to their superpartners, which have fermion number $f+1$ or $f-1$. The fermion number of the time-evolved state $|\psi(t)\rangle=e^{-iHt}|\psi\rangle$ is thus highly constrained, since supersymmetry implies that
@@ III PXP-like fermion model
The sum runs over all sites $i$, and $\langle i\rangle$ denotes the set of sites neighboring site $i$. On a chain, $\langle i\rangle=\{i-1,i+1\}$ subject to the periodic boundary condition $i+N\equiv i$. The annihilation (creation) operators $c_{i}$ ($c_{i}^{\dagger}$) obey the standard fermionic anticommutation relations $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$ and $\{c_{i},c_{j}^{\dagger}\}=\delta_{ij}$. The operators $Q$ and $Q^{\dagger}$ describe the annihilation (creation) of fermions while taking into account the nearest-neighbor blockade constraint, respectively. Because of destructive interference related to the fermionic anticommutation relations, all resulting interactions act locally (nearest-neighbor and next-nearest neighbor). Although the $M_{1}$ model can naturally be defined on any graph, in what follows we restrict the focus to a chain. The resulting supersymmetric Hamiltonian (2) is known as the $M_{1}$ Hamiltonian on the chain and reads
@@
The Hamiltonian consists of the sum of a kinetic part describing constrained hopping and a potential part and we again fix periodic boundary conditions. We obtain the corresponding PXP-like fermion model $H_{\text{PXP}}$ studied in this work by substituting $Q$ as defined in Eq. (6) on the chain geometry in Hamiltonian (4), giving
$H_{\text{PXP}}=\sum_{i=1}^{N}P_{i-1}\big{(}c_{i}^{\dagger}+c_{i}\big{)}P_{i+1}+\mu F\,.$ (8)
This Hamiltonian can naturally be seen as a fermionic equivalent of the PXP model. Indeed, the PXP model is recovered for $\mu=0$ when replacing the fermionic operators $c_{i}^{\dagger}$ and $c_{i}$ by their hard-core bosonic analogs.
@@ IV Periodic revivals
The dynamics of the PXP model show approximate periodic revivals for specific initial states [7, 8]. Periodic revivals for a small set of initial states including, in particular, the $\mathbb{Z}_{2}$-ordered product state, are often seen as a hallmark of quantum many-body scarring [3, 4]. Here, we show analytically that the quench dynamics of the PXP-like fermion model (1) exhibits exact and approximate periodic revivals for properly chosen initial product states.
Exact revivals. Among others, exact revivals can be observed in the dynamics for the $\mathbb{Z}_{2}$-ordered product state. We denote this state as $|\mathbb{Z}_{2}\rangle$, which can be obtained from the vacuum state $|0\rangle$ as $|\mathbb{Z}_{2}\rangle=c_{2}^{\dagger}c_{4}^{\dagger}\dots c_{N}^{\dagger}|0\rangle$, where $N$ is taken to be even. The important observation is that this state is an eigenstate of the $M_{1}$ Hamiltonian as it is annihilated by the kinetic part, having an energy $E=N/2$ from the (diagonal) potential part, where $N$ is the number of sites. Specifically, we have that
@@
where we have used that $Q^{\dagger}|\mathbb{Z}_{2}\rangle=0$ because of the blockade constraint. If we now consider the action of the PXP-like Hamiltonian (1), we find that this state is no longer an eigenstate, but rather gets coupled to its superpartner $Q|\mathbb{Z}_{2}\rangle$, which in turn only couples to $|\mathbb{Z}_{2}\rangle$. This Hamiltonian acts on these two states as
@@
Taking into account the normalization of $Q|\mathbb{Z}_{2}\rangle$, which satisfies $\langle\mathbb{Z}_{2}|Q^{\dagger}Q|\mathbb{Z}_{2}\rangle=(N/2)\langle\mathbb{Z}_{2}|\mathbb{Z}_{2}\rangle=N/2$, the Hamiltonian (1) of the PXP-like fermion model can be restricted to the subspace spanned by the normalized states $\{|\mathbb{Z}_{2}\rangle,Q|\mathbb{Z}_{2}\rangle/\sqrt{N/2}\}$, where it takes the form
@@
$E_{\pm}=\frac{\mu(N-1)}{2}\pm\sqrt{\frac{N}{2}\pm\frac{\mu^{2}}{4}}\,.$ (13)
As such, under the dynamics of Hamiltonian (1), the state $|\mathbb{Z}_{2}\rangle$ will exhibit Rabi oscillations with its superpartner at frequency $\sqrt{2N+\mu^{2}}$. In the absence of a chemical potential, i.e. $\mu=0$, the dynamics is particularly simple, and the fidelity $\mathcal{F}(t)=|\langle\mathbb{Z}_{2}|e^{-iH_{\text{PXP}}t}|\mathbb{Z}_{2}\rangle|^{2}$ can be directly calculated as
@@
The structure discussed here more generally holds for eigenstates of the $M_{1}$ Hamiltonian. For a doublet of eigenstates $|\psi\rangle$ and $Q|\psi\rangle$ where $|\psi\rangle$ has energy $E$ and fermion number $f$, the Hamiltonian (1) takes the form
@@
Approximate revivals. Next to the exact revivals discussed above, approximate revivals can be observed in the dynamics when the system is initialized in a single-fermion product state. The PXP-like fermion model only couples states with a single fermion, $f=1$, to the vacuum state with no fermions and to states with two fermions, i.e. $f=0$ or $f=2$, following our discussion in Sec. II. First note that in the single-fermion basis, the $M_{1}$ Hamiltonian drastically simplifies to a tight-binding Hamiltonian, since in this sector all projector terms evaluate to unity. While the states in the two-fermion basis are more complicated, the relevant eigenstates for the dynamics can be directly obtained from the supersymmetry.
@@
Figure 1: The fidelity $\mathcal{F}(t)$ [Eq. (18)] for the dynamics of the PXP-like fermion model with $N=30$ sites initialized in a single-particle product state. The dashed line shows the envelope $J_{0}^{2}(t/\sqrt{N-2})$ appearing in the large-$N$ approximation as given in Eq. (19). The inset shows the difference $\Delta\mathcal{F}(t)$ between the exact fidelity (18) and the large-$N$ approximation (19). The fidelity exhibits approximate revivals at times scaling as $\sqrt{N}$, and decays as $1/t$ at late times.
@@
where $J_{0}$ is the Bessel function of the first kind. The fidelity exhibits fast oscillations with period $\pi/\sqrt{N-2}$ due to the squared cosine term, and slow oscillations with period $\pi\sqrt{N-2}$ due to the oscillations of the Bessel function. The Bessel function additionally induces an envelope that evolves as $1-t^{2}/[2(N-2)]$ at early times, and as $\sqrt{(N-2)/2}/\,t$ at late times. The fidelity hence exhibits a slow power-law decay as $1/t$. Fig. 1 shows a numerical illustration for a system of size $N=30$. The deviation $\Delta\mathcal{F}(t)$ of approximation (19) from the exact result (18) can be seen to be negligibly small, at least up to moderately late times. We observe from this result that the fidelity of this model exhibits approximate revivals at times scaling with system size as $\sqrt{N}$ for initial single-fermion product states.

### v2

# Periodic revivals from supersymmetry in a fermionic kinetically constrained model

@@ Abstract
Supersymmetry provides a natural playground for the construction of dynamically constrained lattice fermion models. We here illustrate how supersymmetry can be used to construct a fermionic equivalent of the PXP model with an adjustable chemical potential. This model is closely related to the $\mathcal{N}=2$ supersymmetric $M_{1}$ model, inheriting its integrability. The supersymmetric algebra additionally implies that the dynamics exhibit periodic revivals for specific initial states, including the $\mathbb{Z}_{2}$-ordered (every second site occupied) product state. These dynamics are reminiscent to those of the PXP model, a paradigmatic effective model in the field of quantum many-body scars. We draw a further parallel by uncovering eigenstates obeying sub-thermal entanglement scaling at energies given by (plus or minus) square roots of integers and relate these to special eigenstates of the $M_{1}$ model. While we focus on a concrete model, our proposed approach is applicable to more general supersymmetric algebras, where it is expected to lead to non-ergodic dynamics.
@@ I Introduction
Starting from the experimental observation of periodic revivals in the dynamics of programmable atom quantum simulators [1, 2], it is by now well-established that quantum many-body systems with (emergent) constraints can avoid relaxation up to time scales much longer than naively expected [3, 4, 5, 6]. Such dynamics can often be related to a small number of highly non-thermal eigenstates known as quantum many-body scars [7, 8], in loose analogy to the single-particle quantum scars first observed by Heller in 1984 [9, 10, 11]. These eigenstates occur at finite energy density, thereby strongly violating the eigenstate thermalization hypothesis [12, 13, 14].
@@
Several mechanisms leading to long-lived periodic revivals and the emergence of quantum many-body scars have been identified by now. Among these are the presence of an approximate hidden SU(2) algebra [15, 16, 17, 18], a spectrum generating algebra [19, 20, 21, 22, 23, 24], the proximity of the model to integrability [25, 26], or Bethe-ansatz integrability [27, 28]. In this work, we focus on a less-studied mechanism that could induce these phenomena in constrained quantum many-body systems: supersymmetry [29, 30, 31]. As we point out, supersymmetry imposes additional structure on the eigenstates that can not be captured by ordinary conservation laws such as particle number or momentum conservation. For concreteness we consider a model that is additionally integrable and hence non-ergodic, which will allow us to obtain exact results on the eigenstates and the dynamics. However, we emphasize that the presented construction more generally induces non-ergodicity, irrespective of the presence of integrability.
@@
We will consider the supersymmetric algebra underlying the $M_{1}$ model, a relatively well-studied example of a supersymmetric constrained lattice quantum many-body system [32, 33]. It describes short-range interacting spinless fermions on a chain (although it can naturally be defined on any graph), subject to the constraint that two neighboring sites can not be occupied simultaneously. The model is integrable by the Bethe ansatz [34, 35, 32, 33], and can be realized experimentally [36, 37]. Reference [38] recently studied weak ergodicity breaking for the $M_{1}$ model on hypercubic lattice geometries. It was found that a number of highly non-thermal eigenstates, interpreted as quantum many-body scars, can be found thanks to the supersymmetry. In this work, we use the $M_{1}$ model on a chain as a starting point to construct a fermionic equivalent of the (hard-core bosonic) PXP model. The PXP model [39, 40, 7, 8] provides a paradigmatic effective model in studies on quantum many-body scars and weak ergodicity breaking (i.e., the partial breakdown of ergodicity) [3, 4, 5, 6].
@@
The aim of this work is twofold: First, we show how the algebraic structure of supersymmetry can be used to construct kinetically constrained lattice Hamiltonians displaying ergodicity breaking and long-lived periodic revivals. While the procedure holds more generally, we focus on the $M_{1}$ model for concreteness, for which the resulting Hamiltonian represents a fermionic version of the PXP model with an adjustable chemical potential:
@@
expressed in terms of fermionic creation/annihilation operators $c_{i}^{\dagger}$ and $c_{i}$, projectors $P_{i}=1-c_{i}^{\dagger}c_{i}$, and the total fermion number operator $F$. This model is introduced in detail below. We show how the supersymmetry of the $M_{1}$ model leads to exact and approximate periodic revivals in the quench dynamics for specific initial states, including the $\mathbb{Z}_{2}$-ordered (every second site occupied) product state. Periodic revivals for a small set of initial states, typically including the $\mathbb{Z}_{2}$-ordered product state, are often seen as a hallmark of quantum many-body scarring and have also been observed in the dynamics of the PXP model [7, 8] and other models showing weak ergodicity breaking [41].
Putting the Bethe-integrability of the $M_{1}$ model into the mix, second, and of independent interest, we identify special eigenstates of the $M_{1}$ Hamiltonian on a chain. These special eigenstates, occurring at energies given by plus or minus square roots of integers when considering the PXP-like fermion model, support a matrix product state (MPS) description with small bond dimension and a resulting sub-thermal entanglement scaling. Similar special eigenstates at energies given by plus or minus square roots of integers have been observed for the PXP model [42, 43].
@@ II Supersymmetry
Supersymmetric quantum mechanics provides a versatile tool for the construction of constrained lattice fermion models. An $\mathcal{N}=2$ supersymmetric quantum model is described by a Hamiltonian of the form
@@
Observing that $F,H_{S}]=0$ shows that the fermion number is a conserved quantity. We here propose the study of Hamiltonians of the form
@@
where $\mu$ is a free parameter having the interpretation of a chemical potential. Such Hamiltonians have recently appeared in Ref. [44] in the context of central spin models. As one direct illustration of constraints that supersymmetry imposes on the dynamics of this model, we can consider the evolution of the total fermion number. While the total number of fermions $F$ is no longer conserved, the dynamics of the fermion number under the dynamics of $H$ are generically non-ergodic. Consider, for example, a state $|\psi\rangle$ with a fixed fermion number, i.e. $F|\psi\rangle=f|\psi\rangle$. This state is a superposition of eigenstates of the supersymmetric Hamiltonian $H_{S}$ of Eq.  (2) with $f$ fermions. The Hamiltonian $H$ of Eq. (4) only couples these eigenstates to their superpartners, which have fermion number $f+1$ or f1f-1. The fermion number of the time-evolved state $|\psi(t)\rangle=e^{-iHt}|\psi\rangle$ is thus highly constrained, since supersymmetry implies that
@@ III PXP-like fermion model
The sum runs over all sites $i$, and $\langle i\rangle$ denotes the set of sites neighboring site $i$. On a chain, $\langle i\rangle=\{i-1,i+1\}$ subject to the periodic boundary condition $i+N\equiv i$. The annihilation (creation) operators $c_{i}$ ($c_{i}^{\dagger}$) obey the standard fermionic anticommutation relations $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$ and $\{c_{i},c_{j}^{\dagger}\}=\delta_{ij}$. The operators $Q$ and $Q^{\dagger}$ describe the annihilation (creation) of fermions while taking into account the nearest-neighbor blockade constraint, respectively. Because of destructive interference related to the fermionic anticommutation relations, all resulting interactions act locally (nearest-neighbor and next-nearest neighbor). Although the $M_{1}$ model can naturally be defined on any graph, in what follows we restrict the focus to a chain. The resulting supersymmetric Hamiltonian $H_{S}$ as given in Eq. (2) is known as the $M_{1}$ Hamiltonian on the chain and reads
@@
The Hamiltonian consists of the sum of a kinetic part describing constrained hopping and a potential part and we again fix periodic boundary conditions. We obtain the corresponding PXP-like fermion model $H_{\text{PXP}}$ studied in this work by substituting $Q$ as defined in Eq. (6) on the chain geometry in Hamiltonian of Eq. (4), giving the Hamiltonian as given in Eq. (1). This Hamiltonian can naturally be seen as a fermionic equivalent of the PXP model. Indeed, the PXP model is recovered for $\mu=0$ when replacing the fermionic operators $c_{i}^{\dagger}$ and $c_{i}$ by their hard-core bosonic analogs.
@@ IV Periodic revivals
The dynamics of the PXP model show approximate periodic revivals for specific initial states [7, 8]. Periodic revivals for a small set of initial states including, in particular, the $\mathbb{Z}_{2}$-ordered product state, are often seen as a hallmark of quantum many-body scarring [3, 4]. Here, we show analytically that the quench dynamics of the PXP-like fermion model $H_{\text{PXP}}$ as given in Eq. (1) exhibits exact and approximate periodic revivals for properly chosen initial product states. Note however that the integrability of this model, which we will discuss below, precludes the notion of many-body scarring since all eigenstates are non-ergodic.
IV.1 Exact revivals
Among others, exact revivals can be observed in the dynamics for the $\mathbb{Z}_{2}$-ordered product state. We denote this state as $|\mathbb{Z}_{2}\rangle$, which can be obtained from the vacuum state $|0\rangle$ as $|\mathbb{Z}_{2}\rangle=c_{2}^{\dagger}c_{4}^{\dagger}\dots c_{N}^{\dagger}|0\rangle$, where $N$ is taken to be even. The important observation is that this state is an eigenstate of the $M_{1}$ Hamiltonian as it is annihilated by the kinetic part, having an energy $E=N/2$ from the (diagonal) potential part, where $N$ is the number of sites. Specifically, we have that
@@ IV.1 Exact revivals
where we have used that $Q^{\dagger}|\mathbb{Z}_{2}\rangle=0$ because of the blockade constraint. If we now consider the action of the PXP-like Hamiltonian $H_{\text{PXP}}$ of Eq. (1), we find that this state is no longer an eigenstate, but rather gets coupled to its superpartner $Q|\mathbb{Z}_{2}\rangle$, which in turn only couples to $|\mathbb{Z}_{2}\rangle$. This Hamiltonian acts on these two states as
@@
Taking into account the normalization of $Q|\mathbb{Z}_{2}\rangle$, which satisfies $\langle\mathbb{Z}_{2}|Q^{\dagger}Q|\mathbb{Z}_{2}\rangle=(N/2)\langle\mathbb{Z}_{2}|\mathbb{Z}_{2}\rangle=N/2$, the Hamiltonian $H_{\text{PXP}}$ given in Eq. (1) of the PXP-like fermion model can be restricted to the subspace spanned by the normalized states $\{|\mathbb{Z}_{2}\rangle,Q|\mathbb{Z}_{2}\rangle/\sqrt{N/2}\}$, where it takes the form
@@
$E_{\pm}=\frac{\mu(N-1)}{2}\pm\sqrt{\frac{N}{2}+\frac{\mu^{2}}{4}}\,.$ (12)
As such, under the dynamics of Hamiltonian $H_{\text{PXP}}$ of Eq. (1), the state $|\mathbb{Z}_{2}\rangle$ will exhibit Rabi oscillations with its superpartner at frequency $\sqrt{2N+\mu^{2}}$. In the absence of a chemical potential, i.e. $\mu=0$, the dynamics is particularly simple, and the fidelity $\mathcal{F}(t)=|\langle\mathbb{Z}_{2}|e^{-iH_{\text{PXP}}t}|\mathbb{Z}_{2}\rangle|^{2}$ can be directly calculated as
@@
The structure discussed here more generally holds for eigenstates of the $M_{1}$ Hamiltonian. For a doublet of eigenstates $|\psi\rangle$ and $Q|\psi\rangle$ where $|\psi\rangle$ has energy $E$ and fermion number $f$, the Hamiltonian $H_{\text{PXP}}$ as given in Eq. (1) takes the form
@@
IV.2 Approximate revivals
Next to the exact revivals discussed above, approximate revivals can be observed in the dynamics when the system is initialized in a single-fermion product state. The PXP-like fermion model only couples states with a single fermion, $f=1$, to the vacuum state with no fermions and to states with two fermions, i.e. $f=0$ or $f=2$, following our discussion above. First note that in the single-fermion basis, the $M_{1}$ Hamiltonian drastically simplifies to a tight-binding Hamiltonian, since in this sector all projector terms evaluate to unity. While the states in the two-fermion basis are more complicated, the relevant eigenstates for the dynamics can be directly obtained from the supersymmetry.
@@ IV.2 Approximate revivals
Figure 1: The fidelity $\mathcal{F}(t)$ [Eq. (17)] for the dynamics of the PXP-like fermion model with $N=30$ sites initialized in a single-particle product state. The dashed line shows the envelope $J_{0}^{2}(t/\sqrt{N-2})$ appearing in the large-NN approximation as given in Eq. (18). The inset shows the difference $\Delta\mathcal{F}(t)$ between the exact fidelity $\mathcal{F}(t)$ of Eq. (17) and the large-NN approximation as given in Eq. (18). The fidelity exhibits approximate revivals at times scaling as $\sqrt{N}$, and decays as $1/t$ at late times.
@@
where $J_{0}$ is the Bessel function of the first kind. The fidelity exhibits fast oscillations with period $\pi/\sqrt{N-2}$ due to the squared cosine term, and slow oscillations with period $\pi\sqrt{N-2}$ due to the oscillations of the Bessel function. The Bessel function additionally induces an envelope that evolves as $1-t^{2}/[2(N-2)]$ at early times, and as $\sqrt{(N-2)/2}/\,t$ at late times. The fidelity hence exhibits a slow power-law decay as $1/t$. Fig. 1 shows a numerical illustration for a system of size $N=30$. The deviation $\Delta\mathcal{F}(t)$ of the approximation of Eq. (18) from the exact result of Eq. (17) can be seen to be negligibly small, at least up to moderately late times. We observe from this result that the fidelity of this model exhibits approximate revivals at times scaling with system size as $\sqrt{N}$ for initial single-fermion product states.